# Northeast Materials Database (NEMAD): Enabling Discovery of High Transition Temperature Magnetic Compounds

### v1

@@ RESULTS
Random Forest Original 0.80 70K 114K
Balanced 0.83 72K 113K
Ensembled Neural Network Original 0.79 63K 118K
Balanced 0.84 64K 110K
XGBoost Original 0.80 66K 114K
Balanced 0.86 62K 105K
Figure 4: Comprehensive analysis of Curie temperature prediction models using different datasets and evaluation metrics. (a-c) Comparison of predicted vs. actual Curie temperatures for the test set of the Original dataset, derived from the NEMAD database. (d-f) Similar comparison for the Balanced dataset, where some compounds with Curie temperatures below 300K were randomly removed to achieve a balanced distribution of high and low-temperature compounds. For both datasets, the plots show predictions from three models: Random Forest (a,d), Ensemble Neural Network (b,e), and XGBoost (c,f). Each plot displays the coefficient of determination (R2), mean absolute error (MAE), and root mean square error (RMSE) in the top left corner. (g-i) Absolute error distribution and exponential fit across different models for the test set of the Balanced dataset: (g) Random Forest, (h) Ensemble Neural Network, and (i) XGBoost. (j-l) Feature importance plots for Curie temperature prediction models trained on the Balanced dataset, highlighting the top 20 features crucial for accurate predictions: (j) Random Forest, (k) Ensemble Neural Network, and (l) XGBoost. All models utilized features generated from chemical compositions to make predictions. This comprehensive visualization allows for a direct comparison of model performance across different datasets, error distributions, and feature importances, providing insights into the strengths and characteristics of each predictive model in the context of Curie temperature prediction.
@@
Similar method can be developed to predict Néel temperatures of the antiferromagnetic compounds. Details of procedures for data cleaning and model building are explained in the Methods section. We developed and evaluated three distinct models: Random Forest (RF) regressor, Ensemble Neural Network (ENN), and XGBoost. ENN and XGBoost demonstrated superior performance, while RF showed comparatively poorer fit. XGBoost achieved the highest performance with an R2 value of 0.85, followed closely by ENN with an R2 of 0.82. On the other hand, the RF model lagged behind with an R2 of 0.76. The MAE and RMSE further demonstrated the performance of the model. The XGBoost model had the lowest error metrics with an MAE of 32K and RMSE of 67K, while ENN showed a little lower accuracy with an MAE of 36K and RMSE of 73K. However, RF showed substantially higher error rates, with an MAE of 44K and RMSE of 83K.
@@
Figures 5 show the predicted versus actual Néel temperature plot, absolute error plot, and feature importance plot using ENN and XGBoost models. As the figures suggest, the predicted temperatures at the high-temperature region quite deviate from the actual one, this is due to lack of enough training data points in this region. The absolute error plots demonstrate that around 64% data points in the test set have an absolute error of less than 20K for the XGBoost model, whereas in the ENN Model, this proportion is 60%. Feature importance plots show the top 20 most decisive features to predict the Néel temperature of the antiferromagnetic compounds. Both model shows the proportion of Iron(Fe) atom as the most important feature. A significant number of transition metal oxides, including those of iron, manganese, Cobalt, and nickel, exhibit antiferromagnetic behavior. This is due to the superexchange interactions between the magnetic ions mediated by oxygen anions, which often result in anti-parallel alignment of adjacent spins, leading to antiferromagnetic ordering [48, 49]. So the appearance of the oxygen atom on both models’ importance feature list shows that our models somehow captured the pattern correctly.
Figure 5: Evaluation of Ensemble Neural Network and XGBoost models for Néel temperature prediction using features generated from chemical compositions. (a-c) Ensemble Neural Network model results: (a) Predicted vs. actual Néel temperature for the test set, (b) Absolute error distribution, (c) Feature importance plot. (d-f) XGBoost model results: (d) Predicted vs. actual Néel temperature for the test set, (e) Absolute error distribution, (f) Feature importance plot.
@@
Screening High Performance Magnetic Material from Materials Project Database
@@
With the success of training neural network models presented above, we implement our machine learning models in this section on completely unseen data to screen out novel materials with specific properties. To this end, the Materials Project database was selected, from where marked magnetic materials with features like chemical composition, magnetic phase, and stability were downloaded using an API key. First, we removed materials which are already included in our database. Although every material from the Materials Project has a label of its magnetic state, we still implemented our classification model on the dataset to reconfirm the magnetic state, ie, ferromagnetic, antiferromagnetic, or non-magnetic, of each compound. We listed 602 ferromagnetic and 237 antiferromagnetic compounds having the consensus between our classification results and the Materials Project database. We then used our regression models to predict the Curie and Néel temperature of all the above-listed ferromagnetic and antiferromagnetic compounds. Table 4 lists some potential high Curie temperature ferromagnetic compounds with predicted Curie temperatures greater than 500K for at least two best models, XGBoost and ENN, as well as antiferromagnetic compounds with predicted Néel temperatures greater than 100K for at least the XGBoost model. We have detailed all 62 potential high Curie temperature ferromagnetic compounds and 19 antiferromagnetic compounds in the Supplementary Materials. Future experimental verification of these predictions are strongly encouraged.
Table 4: The list of potential high Curie temperature candidates with predicted Curie temperatures and potential AFM compounds with predicted Néel temperature from our three best machine learning models is yet to be verified by experiments. The full list is shown in the Supplementary Materials
Material Composition Predicted Type Ensemble NN (K) XGBoost (K) RF (K)
VFeCoGe FM 609 710 729
FeCo2Ge FM 1051 1068 999
LiFeRh2 FM 636 690 577
AlFeCo2 FM 912 928 907
Ba(FeO2)2 FM 660 659 658
GaFe2Co4Si FM 1042 1013 1007
Fe3RhN FM 677 689 625
Fe3PdN FM 646 673 622
Fe3Co3Si2 FM 1042 939 913
AlFe3C FM 697 784 669
Ga3Fe4Co8Si FM 1151 1157 1010
Li2Fe12P7 FM 734 655 643
MnCo2Si FM 823 991 662
MnCo2Sb FM 617 691 672
MnCo2Sn FM 802 798 767
NbAlCo2 FM 770 609 605
MnGaCo2 FM 740 668 648
Mn2GaCo4Ge FM 804 722 658
MnCo2Ge FM 767 900 803
TaAlCo2 FM 729 609 516
MnAlCo2 FM 805 757 700
Sr6(CoO3)5 FM 642 746 620
TiFeO3 AFM 60 137
Sr2FeBrO3 AFM 139 300
TaFeO4 AFM 53 122
BaFe2P2O7F2 AFM 132 132
Cr3(FeO6)2 AFM 136 123
Lu(Al2Fe)4 AFM 99 125
Sr2Co(BrO)2 AFM 63 116
Ba2Mn2Se2OF2 AFM 105 102
La2Mn2Se2O3 AFM 150 151
@@ Discussion
It is interesting to note that our database only includes 35% antiferromagnetic materials, far less than ferromagnets. One reason may be because of the presence of many ferromagnetic alloys in the database. The number of alloys is in principle unbounded and those containing magnetic elements such as iron are more likely to be ferromagnetic due to their itinerancy nature. On the other hand, there exist many antiferromagnets in transition metal oxides. Experimental reports for this community were published in journals outside of Elsevier. It is thus important to expand the coverage of articles to include other publishers such as the American Physical Society. A large database on magnetic materials is expected.
@@
Our database also include the structural information of magnetic materials. This can be used in building a more powerful models, such as the graph neural network, in predicting magnetic properties. We also built a simple XGBoost machine learning model to incorporate these structure features like crystal system and space group (detailed included in the Supplementary Materials). The XGBoost model achieved an R2 value of 0.86, MAE of 52K, and a RMSE of 89K. The MAE and RMSE values are slightly improved compared to the model without structural information.
@@ METHODS
Figure 6 illustrates the detailed workflow of the method used in this work. We began by compiling a comprehensive list of $40,000$ Digital Object Identifiers (DOIs) of scientific articles related to Magnetic Materials by searching relevant keywords (Ferromagnetic, Anti-ferromagnetic, Curie, and Néel) and deploying web scraping tool in various journal websites. We downloaded the article for all listed DOIs using the authenticated API request. These articles were primarily obtained in XML format, which often contains a hierarchical structure with nested elements. We developed a custom XML parsing script to extract the full content of articles, including tabular data, from pre-downloaded XML-formatted documents. The script converts all unstructured textual and tabular information into plain text format and then saves it as a CSV file with the following features.
@@
GPTArticleExtractor[47] method provides a way to make a comprehensive material property database by automatically extracting data from scientific articles with high precision and recalls. This method leverages the cutting edge Large Language Models like GPT-3.5 by incorporating the OpenAI API key. One can extract their desired information from articles by just modifying the prompts used in the workflow. So effective prompts are crucial for the accuracy and efficiency of this method.
For this study, we listed the desired properties to extract from articles, including material chemical composition, magnetic phase transition temperatures (Curie, Néel, and Curie-Weiss), structural information (Crystal Structure, Lattice Structure, Lattice Parameters, and Space Group), and magnetic properties (Coercivity, Magnetization, Magnetic Moment, Remanence, and Susceptibility). We carefully reviewed literature from a range of publications to comprehend the trends and writing styles associated with these attributes to create effective prompts. We formulated the question prompts to extract the custom need information. For instance, to extract material chemical composition, we formulated the prompt like this " Provide an exact chemical composition of the magnetic materials studied in the article, some materials’ composition are expressed with the variable proportion of elements. So include any value of the variable mentioned."
Large Language Models are not still free from the limitation of the input tokens. It is less effective if we use the entire sections of the paper as an input of these models. To include only information related to the question prompt in the input, GPTArticleExtractor tokenized the articles by segmenting the text into chunks of 500 tokens each. Following that, these segments were examined in a vector space to determine word similarity using metrics like Euclidean distance. By comparing the vectorized text segments with particular question prompts, the model figured out the five most relevant segments using Facebook AI Similarity Search (FAISS) [50]. This method allowed us to focus on the most relevant information, improving answer quality while minimizing input tokens and computing expenses. After selecting the most correlated information from the articles, it generates the output with desired properties by utilizing the summarization prompt. To ensure consistency in output results, we employed one-shot prompting where the prompt included an example output format. This technique enabled the generation of accurate and structured output data.
Although this workflow performs well for handling longer articles with the help of GPT 3.5 and GPT 4 models, it still has certain shortcomings. The relevant segment parts occasionally missed important materials and properties information when multiple entries were present in a single article and struggled to extract stoichiometric ratios presented in variable forms. We modified this workflow with the advent of the powerful GPT-4o model. We included all the details of the article except the introduction and reference sections as input in the modified workflow. This all-inclusive strategy performed better than the prior technique, producing outputs that were more comprehensive, precise, and organized. This approach managed various material compositions and properties present in a single article by accurately creating lists of dictionaries, where each entry includes all the information associated with a single chemical composition. Some chemical compositions extracted by our model missed certain properties. Therefore, we established a rule for including these materials in the database: a chemical composition must contain at least phase transition temperature information to be included. Finally, We built the automated database of magnetic materials with 26,706 entries.
@@
Average Atomic $\text{Average Atomic Number}(\bar{Z})=\sum_{i=1}^{84}p_{i}Z_{i}$ (3)

### v2

@@ Results
Random Forest Original 0.82 71K 115K
Balanced (Undersampled) 0.85 65K 104K
Ensembled Neural Network Original 0.84 60K 110K
Balanced (Undersampled) 0.85 56K 103K
XGBoost Original 0.84 62K 107K
Balanced (Undersampled) 0.87 56K 97K
[…]
To further assess the model’s predictive accuracy, we conducted an error analysis using the absolute error plots. Figures 4g, 4h, and 4i presents these plots for all three models across balanced dataset. We fitted an exponential curve to each error distribution to characterize the error pattern. In XGBoost and ENN, 70% of the test data have an absolute error of less than 50K, whereas in RF model, this proportion is 62%.
Figure 4: Evaluation of Curie temperature prediction models on original and stratified-balanced datasets. (a–c) Predicted versus actual Curie temperatures for the test set of the original dataset using three models: Random Forest (a), Ensemble Neural Network (b), and eXtreme Gradient Boosting (XGBoost) (c). (d–f) Corresponding predictions on a balanced dataset created using stratified undersampling. To construct this dataset, the Curie temperature range was divided into bins, and samples from the overrepresented low-temperature region were undersampled to create a more uniform temperature distribution. Each plot reports the coefficient of determination (R2), mean absolute error (MAE), and root mean squared error (RMSE), and includes confidence intervals based on ensemble standard deviation. (g–i) Absolute error distributions and fitted exponential curves for the balanced dataset across the three models: Random Forest (g), Ensemble Neural Network (h), and XGBoost (i). (j–l) Feature importance plots showing the top 20 most influential features for models trained on the balanced dataset: Random Forest (j), Ensemble Neural Network (k), and XGBoost (l). All models were trained on features derived from the chemical composition of materials in the NEMAD database. The figure provides a comparative assessment of model performance, uncertainty, and feature relevance under both original and balanced training conditions.
To better understand the contribution of each feature in predicting curie temperature, we computed model-specific feature importances using methods appropriate for each algorithm. For the RF model, feature importance was calculated as the average reduction in impurity across trees in the ensemble. For the XGBoost model, we used the average gain metric, which quantifies the improvement in the loss function brought by each feature across all decision trees. For the ENN, permutation importance was applied, which measures the increase in prediction error when the values of a given feature are randomly shuffled. In all cases, feature importances were averaged across 30 independently trained models, and standard deviations were reported to reflect the variability across the ensemble. Figures 4j, 4k, and 4l display the top 20 features ranked by importance for each model. Notably, features such as the proportion of Fe and Co atoms, the proportion of high-Curie-temperature elements, the average magnetic moment, and the presence of rare-earth elements consistently appeared among the most influential predictors across all models.
@@
Similar methods can be developed to predict Néel temperatures of antiferromagnetic compounds. Details of procedures for data cleaning and model building are explained in the Methods section. We trained and evaluated three distinct models: RF regressor, ENN, and XGBoost. Among these, ENN and XGBoost demonstrated consistently strong performance, while RF showed slightly lower accuracy in error metrics. XGBoost achieved an R2 value of 0.83, followed by ENN with an R2 of 0.80. The RF model yielded a comparable R2 value of 0.81.
The MAE and RMSE further illustrate the performance of the models. The XGBoost model had an MAE of 38 K and RMSE of 72 K, while ENN achieved an MAE of 38 K and RMSE of 76 K. The RF model showed slightly higher error rates, with an MAE of 43 K and RMSE of 76 K. Although the differences in performance are modest, XGBoost and ENN show consistently competitive results across both metrics.
Figure 5 presents the predicted versus actual Néel temperature plots, along with absolute error distributions and feature importance plot for the RF, ENN, and XGBoost models. Similar to the case of Curie temperature prediction, the deviation in predicted values is more pronounced at higher temperatures, likely due to the limited availability of high-Néel-temperature training samples. The absolute error plots demonstrate that approximately 64% of data points in the test set have an absolute error below 25 K for the XGBoost model, whereas in the ENN model, this proportion is 63%. For the Random Forest model, around 60% of predictions fall within this error threshold.
Figure 5: Evaluation of machine learning models for Néel temperature prediction using stratified undersampling. Top row (a–c): Predicted versus actual Néel temperatures on the test set using three models-Random Forest (a), Ensemble Neural Network (b), and eXtreme Gradient Boosting (c)-trained on a balanced dataset created via stratified undersampling. Each plot displays the coefficient of determination (R2), mean absolute error (MAE), and root mean squared error (RMSE). The predictions are color-coded by the standard deviation across ensemble predictions, and error bars reflect the confidence interval associated with each prediction. Middle row (d–f): Distribution of absolute prediction errors on the test set for the same models, with an exponential fit overlaid and the mean absolute error marked by a dashed red line. Bottom row (g–i): Top-20 feature importance plots for each model, showing the most influential descriptors used in predicting the Néel temperature.
Each scatter plot includes both error bars and color-coded points showing the standard deviation of predictions across ensemble models. This provides a visual indication of prediction uncertainty for all three models. Compared to RF and XGBoost, the ENN shows noticeably larger confidence intervals. This reflects higher uncertainty in its predictions. The ENN is a highly nonlinear model, which tends to be more sensitive to small changes in data. In our case, we used stratified undersampling to train 30 separate ENN models, each on a different balanced subset of the data. The undersampling was applied only to the low Néel temperature region, meaning that each model was trained on a different subset of low-temperature samples. This introduces variation in model behavior and results in higher standard deviation across predictions. These wider confidence intervals suggest that the ENN predictions are less reliable for compounds with lower Néel temperatures. In contrast, the RF and XGBoost models show narrower confidence intervals, indicating more stable and consistent predictions across ensemble members.
@@
Feature importance plots show the top 20 most influential features in predicting the Néel temperature of antiferromagnetic compounds. All three models identified the proportion of iron atom (Fe) in the chemical composition as an important feature. A significant number of transition metal oxides, including those of iron, manganese, cobalt, and nickel, exhibit antiferromagnetic behavior. This arises from superexchange interactions between magnetic ions mediated by oxygen anions, which favor anti-parallel spin alignment, leading to antiferromagnetic ordering [51, 52]. The consistent appearance of oxygen among the top-ranked features in all models further supports that the models have learned physically meaningful patterns.
@@
Screening High Performance Magnetic Material from External Databases
@@
Building on the successful training and evaluation of our machine learning models, we now apply them to a separate, previously unseen dataset to identify materials with targeted properties. To this end, the Materials Project database was selected, from where marked magnetic materials with features like chemical composition, magnetic phase, and stability were downloaded using an API key. First, we removed materials which are already included in our database. Although every material from the Materials Project has a label of its magnetic state, we still implemented our classification model on the dataset to reconfirm the magnetic state, i.e., ferromagnetic, antiferromagnetic, or non-magnetic, of each compound. We listed 602 ferromagnetic and 237 antiferromagnetic compounds having the consensus between our classification results and the Materials Project database. We then used our regression models to predict the Curie and Néel temperature of all the above-listed ferromagnetic and antiferromagnetic compounds.
In addition to the Materials Project, we also applied our models to another source of magnetic materials: a curated set of thermodynamically stable Heusler and half-Heusler compounds reported in a high-throughput DFT study [53]. These compounds were first checked to remove any duplicates already present in our database. Then, we applied the same classification and regression models to determine their magnetic state and predict their Curie or Néel temperatures.
Table LABEL:tab:Table4 presents a selection of 25 newly identified ferromagnetic compounds with predicted Curie temperatures exceeding 500K in at least two models, along with 13 antiferromagnetic compounds with predicted Néel temperatures above 100K based on the XGBoost model. These candidates were drawn from both the Materials Project and the DFT-verified Heusler datasets. In total, our screening identified 32 high-probability magnetic compounds, of which 7 have since been found in the literature with experimentally reported Curie temperatures, validating our model predictions. The remaining 25 represent candidates not previously reported with experimental magnetic ordering. Future experimental verification of these predictions is strongly encouraged.
Table 4: List of candidate ferromagnetic and antiferromagnetic compounds with high predicted Curie and Néel temperatures, as estimated by our three best-performing machine learning models. Compounds marked with * have experimental ordering temperatures reported in the literature (see Comment column), while all other candidates remain to be experimentally verified.
Composition Type Crystal System XGBoost (TC) ENN (TC) RF (TC) Comment
VFeCoGe FM Cubic $500.0\pm 57.2\text{K}$ $256.0\pm 142.4\text{K}$ $556.0\pm 19.0\text{K}$
LiNbFeO4 FM Tetragonal $511.7\pm 78.0\text{K}$ $587.9\pm 202.2\text{K}$ $424.1\pm 21.4\text{K}$
Fe2NiP FM Tetragonal $582.1\pm 36.1\text{K}$ $381.7\pm 83.1\text{K}$ $583.1\pm 10.8\text{K}$
GaFe2Co4Si FM Trigonal $1004.9\pm 35.7\text{K}$ $1009.7\pm 60.8\text{K}$ $978.8\pm 16.8\text{K}$
Fe3PdN FM Cubic $510.7\pm 54.7\text{K}$ $600.1\pm 174.9\text{K}$ $478.1\pm 27.3\text{K}$
AlFe3H FM Cubic $569.0\pm 61.8\text{K}$ $711.0\pm 101.9\text{K}$ $550.5\pm 36.3\text{K}$
Fe3Rh FM Cubic $645.6\pm 39.6\text{K}$ $878.6\pm 125.9\text{K}$ $572.8\pm 20.9\text{K}$
Hf(GaFe)6 FM Orthorhombic $527.0\pm 31.4\text{K}$ $500.0\pm 108.3\text{K}$ $466.1\pm 12.8\text{K}$
Fe3Co3Si2 FM Trigonal $45.3\pm 41.5\text{K} 99$$99.0\pm 68.8\text{K}$ $900.9\pm 20.7\text{K}$
AlFe3C FM Cubic $566.8\pm 58.3\text{K}$ $388.3\pm 188.1\text{K}$ $502.3\pm 40.0\text{K}$
Ga3Fe4Co8Si FM Trigonal $998.7\pm 36.3\text{K}$ $1046.2\pm 72.0\text{K}$ $938.9\pm 13.9\text{K}$
Zr(GaFe)6 FM Orthorhombic $500.6\pm 34.1\text{K}$ $520.6\pm 168.7\text{K}$ $482.9\pm 14.1\text{K}$
Mn2GaCo4Ge FM Trigonal $658.0\pm 43.6\text{K}$ $613.9\pm 148.3\text{K}$ $455.8\pm 38.2\text{K}$
Mn4Cr(Co2Ge)5 FM Trigonal $557.3\pm 80.4\text{K}$ $661.8\pm 183.5\text{K}$ $444.9\pm 48.6\text{K}$
Fe3RhN FM Cubic $524.0\pm 54.2\text{K}$ $677.6\pm 170.1\text{K}$ $472.3\pm 26.6\text{K}$
Fe4Li3SbO8 FM Monoclinic $712.6\pm 53.5\text{K}$ $691.0\pm 189.7\text{K}$ $630.9\pm 16.2\text{K}$
MnZn3(CrSe2)8 FM Trigonal $788.3\pm 117.4\text{K}$ $982.5\pm 115.1\text{K}$ $560.5\pm 75.4\text{K}$
VFe2BO5 FM Orthorhombic $552.7\pm 70.4\text{K}$ $288.8\pm 144.9\text{K}$ $514.8\pm 18.2\text{K}$
GaFeNiCo FM Cubic $923.2\pm 58.8\text{K}$ $848.4\pm 151.5\text{K}$ $872.4\pm 43.1\text{K}$
RhCoGaFe FM Cubic $60.8\pm 60.4\text{K} 66$$60.8\pm 60.4\text{K}$ $609.0\pm 18.1\text{K}$
FeGaRuCo FM Cubic $644.2\pm 56.2\text{K}$ $617.5\pm 137.4\text{K}$ $595.2\pm 20.3\text{K}$
GeFeNiCo FM Cubic $740.7\pm 74.0\text{K}$ $856.1\pm 187.1\text{K}$ $633.8\pm 28.6\text{K}$
TcCoGeFe FM Cubic $637.0\pm 56.1\text{K}$ $692.3\pm 161.1\text{K}$ $567.5\pm 19.0\text{K}$
Fe2RhGa FM Cubic $569.7\pm 43.2\text{K}$ $595.9\pm 147.6\text{K}$ $483.8\pm 15.1\text{K}$
Co2NiGe FM Tetragonal $664.4\pm 70.8\text{K}$ $846.1\pm 184.2\text{K}$ $483.6\pm 24.0\text{K}$
Fe3PtN FM Cubic $614.4\pm 43.1\text{K}$ $706.0\pm 144.5\text{K}$ $445.2\pm 17.3\text{K}$ 642K [62]
Y(GaFe)6 FM Orthorhombic $463.7\pm 20.1\text{K}$ $431.7\pm 52.1\text{K}$ $437.3\pm 17.8\text{K}$ 500 K [63]
Co2MnSb FM Cubic $590.7\pm 34.3\text{K}$ $592.9\pm 110.2\text{K}$ $533.1\pm 41.9\text{K}$ $600\pm 10\text{K}$ [64]
Fe2CuGa FM Tetragonal 690.9 ± 73.3 K 716.2 ± 173.6 K 565.9 ± 16.8 K 798 K [65]
GeCrCoFe FM Cubic $634.1\pm 85.7\text{K}$ $416.4\pm 208.5\text{K}$ $623.9\pm 30.1\text{K}$ 866 K [66]
NbAlCo2 FM Cubic $329.2\pm 62.2\text{K}$ $386.6\pm 175.2\text{K}$ $382.8\pm 15.7\text{K}$ 383 K[67]
Fe2RuGe FM Cubic $573.2\pm 27.4\text{K}$ $571.4\pm 121.4\text{K}$ $448.6\pm 25.3\text{K}$ 860 K [68]
Sr2FeBrO3 AFM Tetragonal $284.8\pm 30.6\text{K}$ $119.1\pm 65.8\text{K}$ $249.5\pm 6.8\text{K}$
Sr2FeClO3 AFM Tetragonal $290.2\pm 34.5\text{K}$ $157.6\pm 77.0\text{K}$ $238.5\pm 8.6\text{K}$
BaFeP(O2F)2 AFM Monoclinic $224.1\pm 40.9\text{K}$ $203.3\pm 163.8\text{K}$ $128.0\pm 9.7\text{K}$
K2NaFeO3 AFM Orthorhombic $137.6\pm 43.7\text{K}$ $81.0\pm 101.0\text{K}$ $160.8\pm 10.0\text{K}$
Li2FeO3 AFM Monoclinic $138.3\pm 31.1\text{K}$ $34.3\pm 18.8\text{K}$ $138.3\pm 11.4\text{K}$
Ba3Fe2Br2O5 AFM Cubic $189.5\pm 49.5\text{K}$ $54.0\pm 52.4\text{K}$ $181.1\pm 14.6\text{K}$
Cr3(FeO6)2 AFM Monoclinic $101.0\pm 32.9\text{K}$ $95.4\pm 66.8\text{K}$ $69.5\pm 6.9\text{K}$
LiFeAsO4 AFM Orthorhombic $110.1\pm 30.4\text{K}$ $32.9\pm 11.3\text{K}$ $97.8\pm 10.0\text{K}$
NaFeAsO4F AFM Monoclinic $138.7\pm 39.8\text{K}$ $56.3\pm 52.1\text{K}$ $91.7\pm 6.6\text{K}$
SrCrHO2 AFM Tetragonal $105.4\pm 30.4\text{K}$ $127.5\pm 71.7\text{K}$ $85.5\pm 6.0\text{K}$
IrFeSnHf AFM Cubic $129.3\pm 29.4\text{K}$ $160.8\pm 132.0\text{K}$ $148.6\pm 9.9\text{K}$
HfSnRhFe AFM Cubic $32.2\pm 26.6\text{K} 11$$13.6\pm 81.5\text{K}$ $152.1\pm 9.2\text{K}$
IrFeSnTi AFM Cubic $157.1\pm 32.6\text{K}$ $313.1\pm 179.7\text{K}$ $168.6\pm 8.6\text{K}$
@@ Discussion
It is interesting to note that our database only includes 30% antiferromagnetic materials, far less than ferromagnets. One reason may be because of the presence of many ferromagnetic alloys in the database. The number of alloys is in principle unbounded and those containing magnetic elements such as iron are more likely to be ferromagnetic due to their itinerancy nature. On the other hand, there exist many antiferromagnets in transition metal oxides. Experimental reports for this community were published in journals outside of Elsevier and APS. It is thus important to expand the coverage of articles to include other publishers such as the Springer. A large database on magnetic materials is expected.
@@
Our database also include the structural information of magnetic materials. This can be used in building a more powerful models, such as the graph neural network, in predicting magnetic properties. We also built a simple XGBoost machine learning model to incorporate these structure features like crystal system and space group (detailed included in the Supplementary Information). The XGBoost model achieved an R2 value of 0.83, MAE of 52K, and a RMSE of 98K. The MAE and RMSE values are slightly improved compared to the model without structural information.
@@ Methods
Figure 6 illustrates the detailed workflow of the method used in this work. We began by compiling a comprehensive list of $100,000$ Digital Object Identifiers (DOIs) of scientific articles related to Magnetic Materials by searching relevant keywords (Ferromagnetic, Anti-ferromagnetic, Curie, and Néel, Coercivity, Magnetization, Magnetocrystalline Anisotropy, Remanence) from journal platforms maintained by Elsevier and the American Physical Society. We downloaded the article for all listed DOIs using the authenticated API request. These articles were primarily obtained in XML format, which often contains a hierarchical structure with nested elements. We developed a custom XML parsing script to extract the full content of articles, including tabular data, from pre-downloaded XML-formatted documents. The script converts all unstructured textual and tabular information into plain text markdown format and then saves it as a CSV file with the following features.
GPTArticleExtractor[50] method provides a way to make a comprehensive material property database by automatically extracting data from scientific articles. This method leverages the cutting edge Large Language Models like GPT-3.5 by incorporating the OpenAI API key. One can extract their desired information from articles by just modifying the prompts used in the workflow. So effective prompts are crucial for the accuracy and efficiency of this method.
Although the original workflow demonstrated strong performance when processing lengthy articles using GPT-3.5 and GPT-4 models, it exhibited several limitations. First, it was unable to extract information from tables, which often contain key material properties not explicitly mentioned in the body text. This omission increased the risk of missing essential data. Second, the prompts used in the previous workflow lacked depth and specificity. In several instances, the language models failed to interpret the prompt correctly, particularly when articles reported multiple compositions and their corresponding properties. As a result, some relevant compositions were either omitted or extracted in inconsistent formats, making downstream processing and integration into the database challenging. Moreover, when chemical compositions were expressed in variable stoichiometric forms, such as FexCo1-x, the previous workflow often failed to extract the actual values of the variables. These values are crucial in materials science, and their omission renders the extracted composition scientifically meaningless. Another significant limitation was the workflow’s dependence solely on API-based access to XML-formatted articles. It was not designed to process Portable Document Format (PDF) articles, scanned documents, or older scientific handbooks from the 20th century that are typically stored as image-based files.
@@
To overcome these limitations, we modified the previous GPTArticleExtractor workflow by leveraging the capabilities of the more powerful GPT-4o model. The enhanced pipeline supports XML articles accessed via API, as well as PDF documents, scanned image-based articles, and digitized historical handbooks. By expanding the input modalities and refining the prompt design, this upgraded workflow provides a more robust and inclusive data extraction process for constructing a comprehensive materials property database.
We listed the desired properties to extract from articles, including material chemical composition, magnetic phase transition temperatures (Curie, Néel, and Curie-Weiss), structural information (Crystal Structure, Lattice Structure, Lattice Parameters, and Space Group), and magnetic properties (Coercivity, Magnetization, Magnetic Moment, Remanence, and Susceptibility). We carefully reviewed literature from a range of publications to comprehend the trends and writing styles associated with these attributes to create effective prompts. For each property group, we designed a corresponding question prompt tailored to extract that specific type of information from the article. For instance, to extract material chemical composition, we formulated the prompt like this: “Provide an exact chemical composition of the magnetic materials studied in the article. Some materials’ compositions are expressed with variable proportions of elements, so include any values of such variables if mentioned." Similar prompts were constructed for structural information, magnetic transitions, and magnetic properties, forming a modular querying framework.
Large Language Models are still constrained by input token limitations. It is less effective to use entire sections of a paper as input to these models, especially in the case of longer articles where non-relevant content may dilute the model’s focus and exceed token limits. To address this, we designed a two-stage approach where question prompt was used to retrieve only the most relevant content from long-form articles. We first tokenized the articles by segmenting the text into chunks of 500 tokens each. Following that, these segments were examined in a vector space to determine word similarity using metrics like Euclidean distance. By comparing the vectorized text segments with particular question prompts, the model figured out the five most relevant segments using Facebook AI Similarity Search (FAISS) [54]. This targeted input strategy allowed the model to focus on the most contextually relevant text, thereby improving response accuracy while reducing input token count and computing expenses. After identifying and collecting the most relevant text segments from the articles, the LLM was used to extract the final structured information using a summarization prompt. To ensure output consistency and format adherence, we used a one-shot prompting strategy in which the prompt also contained a single illustrative example of the expected output format. This technique enabled the generation of accurate and structured outputs. See the supplementary information for the full prompt templates.
Even with the use of a more powerful model and improved prompts, some important information was still missed. In particular, the model struggled with chemical compositions that had variable stoichiometry. This issue often occurred when an article contained many different compositions and discussed them throughout the entire article. The tokenized chunks sometimes left out useful parts of the text, which led to incomplete data extraction. While the updated workflow was designed to manage token limitations in earlier models, recent advances in language models, such as GPT-4o, have increased input token capacity. As a result, it is now possible to process nearly the full content of an article. We found that removing non-informative section such as references allowed us to fit the majority of the relevant content within the model’s token limit. This adjustment further improved extraction accuracy, as it preserved the core technical sections containing material property data. We found that this comprehensive approach outperformed the previous technique, resulting in outputs that were more complete, accurate, and well-structured. This approach managed various material compositions and properties present in a single article by accurately creating list of jsons, where each entry includes all the information associated with a single chemical composition. Some outputs extracted by our model included only the chemical composition but lacked all other property information. Therefore, we established an inclusion criterion for the database: a material entry must contain at least one phase transition temperature or at least one magnetic property in addition to its composition.
To handle the PDF articles, we utilized a PDF parser [55, 56] to first convert the PDF into a structured markdown format. Although it is technically possible to pass the raw PDF directly to the language model, we observed that this often leads to suboptimal extraction performance. This is primarily because raw PDFs lack a consistent textual structure and require an intermediate parsing step to resolve layout complexities such as multi-column formatting, table structures, and chemical notations. During PDF to text conversion, crucial associations between material properties and compound names can become fragmented due to token chunking and the loss of layout context, which degrades extraction model accuracy. In contrast, the markdown format preserves hierarchical and semantic cues (e.g., headings, bullet points, tables) that significantly improve the model’s ability to extract and organize information reliably. We then applied the same procedure to process all PDF articles and consolidated the extracted content into a single CSV file, where each row corresponds to a distinct article in markdown form. This structured dataset, along with carefully designed prompts, was subsequently passed to the language model to generate the required information in a consistent and organized format.
Likewise, to extract information from scanned PDF-formatted scientific articles, we used the Gemini 2.0 Flash model to first convert them into a structured markdown format. This model employs advanced multimodal capabilities to process both image and text inputs. Since scanned PDFs often contain only rasterized page images without any embedded text layer, direct information extraction is not feasible. Therefore, the model first applies optical character recognition (OCR) to each page to recover the underlying textual content. Gemini 2.0 Flash leverages Google’s proprietary OCR and layout detection tools to accurately extract not only the raw text but also its structural context-such as headings, paragraph blocks, tables, and figure captions. Following OCR, the model reconstructs the document’s logical hierarchy and semantic structure, mapping it into markdown syntax with appropriate formatting cues (e.g., headings, bullet points, and table structures). This conversion is essential for improving the language model’s ability to identify and extract relevant scientific information consistently. While direct conversion is feasible for shorter scientific articles, it becomes impractical for longer documents such as handbooks due to the input token limitations of the language model. To address this, we first segmented each handbook into individual pages and processed each page separately through the model. The output from each page was then stored in markdown format within a consolidated CSV file. We applied the same structured prompting and extraction workflow to these segments to ensure consistency and completeness of the final output.
This revised workflow is designed to handle diverse document types, including API-accessed articles, standard and scanned PDFs, and historical handbooks. It enables reliable extraction of scientific content into structured, machine-readable formats. Finally, using this pipeline, we built an automated database of magnetic materials comprising 67,573 entries.
@@
$\text{Average Atomic Number}(\bar{Z})=\sum_{i=1}^{84}p_{i}Z_{i}$ (3)